# Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems

### v1

@@ Appendix B Minecraft Comprehensive Skill Library
Our proposed skill library is implemented based on MineflayerPrismarineJS (2023).To simulate the process of solving prerequisite steps in the actual game to eventually obtain specific items, this skill library has introduced a recursive task decomposition mechanism. From the implementation level, all interfaces can recursively invoke each other or themselves and pass parameters.
B.1 Basic Skill Interfaces
  • obtainItem(bot, count, type): Automates the collection of items.
  • mineItem(bot, count, type, explore_direction, explore_time): Mines using specific tools. If the exploration parameters are not specified, the exploration direction is chosen randomly. If the item type is a subterranean mineral like diamond, the exploration direction is set to down (0, -1, 0). The exploration time limit defaults to five minutes (6000 ticks in Minecraft).
  • craftItem(bot, count, type, need_crafting_table): Crafts items. If no parameter is specified, the crafting is done using a crafting table by default.
  • smeltItem(bot, count, type, fuel): Smelts or cooks items in a furnace. If no parameters are provided, coal is used as the default fuel.
  • collectItem(bot, count, type, function): Collects items by killing animals or using other special methods.
  • chatMessage(bot, message, team_name): Sends a chat message to the team.
  • getItemFromChest(bot, chest_position, items_to_get): Retrieves specific items from a chest at a given location.
  • depositItemIntoChest(bot, chest_position, items_to_deposit): Deposits specific items into a chest at a given location.
  • combatWithEntity(bot, mob_name, weapon, loop): Automatically equips the highest quality armor and weapon from the inventory to fight with an entity (e.g., animals or hostile mobs). If no parameters are specified, the default weapon is a sword. The loop parameter determines whether to continue fighting the same type of entity.
  • combatWithPlayer(bot, player_name, weapon): Automatically equips the highest quality armor and weapon from the inventory to fight with a specific player. If no parameters are provided, the default weapon is a sword.
  • initialInventory(bot, item_dict): Initializes the player’s inventory with specified items.
  • equipBestToolOrArmor(bot, type): Automatically equips the highest quality tool, weapon, or armor of the specified type.
  • listenChat(bot, player_name): Continuously listens to a player’s chat messages.
B.2 Recursive-related Interfaces
  • preTool(item): Retrieves the minimum prerequisite tools required to gather a specific item.
@@ B.2 Recursive-related Interfaces
  • preItem(item): Retrieves the prerequisite items required to craft a specific item and indicates whether a crafting table is necessary.
  • preSmelt(item): Retrieves the prerequisite items required to smelt a specific item.
  • preCollect(item): Retrieves the prerequisite entities required to collect a specific item.
  • getFunc(item): Retrieves the method used to collect a specific item.
This library is scalable and can be easily extended to accommodate new items, recipes, and tasks. Currently, it supports the collection of over 790 types of items across all three dimensions of Minecraft (Overworld, Nether, and End). However, the library does not yet support some specialized gameplay features. Future updates will aim to expand the library’s functionality, adding support for more complex tasks.
@@ Appendix C Experiment Details
Experiments were conducted using the Qwen-Plus model API provided by Alibaba Cloud. In the multiplayer local area network (LAN) server of Minecraft, agents connected to the game via different ports and interacted with the environment and other players.
C.1 Resource Collection Task
We implemented a comprehensive Minecraft resource collection benchmark, which allows multi-agent collaboration to automatically collect resources based on a given requirement dictionary, making this benchmark general and scalable. We conducted experiments on the following representative resource collection tasks:
Task Definitions:
  • Iron Tool Set: { ’iron_pickaxe’: 1, ’iron_shovel’: 1, ’iron_hoe’: 1, ’iron_axe’: 1 } — A set of commonly used iron tools in Minecraft.
  • Diamond Armor: { ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1 } — A full set of diamond armor in Minecraft.
  • Redstone Devices: { ’repeater’: 1, ’piston’: 1, ’dropper’: 1 } — Common redstone components and devices in Minecraft.
  • Navigation Kit: { ’compass’: 1, ’clock’: 1, ’map’: 1 } — A set of tools commonly used for navigation in Minecraft.
  • Transport System: { ’minecart’: 1, ’rail’: 16, ’powered_rail’: 6 } — Minecarts and tracks used to complete a transportation system.
  • Food Supplies: { ’beef’: 1, ’chicken’: 1, ’porkchop’: 1 } — Common animal meat food items in Minecraft.
  • Building Materials: { ’stone_bricks’: 4, ’glass’: 4, ’iron_door’: 1 } — Building materials commonly used for constructing doors, windows, and walls.
  • Storage System: { ’hopper’: 1, ’chest’: 1, ’barrel’: 1 } — Tools commonly used for storing items in Minecraft.
@@ C.1 Resource Collection Task
C.2 Boss Combat Task
@@
We manually implemented three representative and challenging task scenarios: defeating the Elder Guardian in an ocean monument in the Overworld, defeating the Wither in the Nether, and defeating the Ender Dragon in the End. Defeating the Ender Dragon represents the final achievement in the game. Prior to each combat, agents were equipped with standardized combat resources, including a full set of diamond armor (validated as efficiently collectible through our resource collection experiments), bow and arrows, and some consumables. In all experiments, the standardized combat resources remained consistent. A single LLM call initialized resource allocation, distributing equipment to each agent through a dictionary-based assignment system. Agents were then teleported to designated combat locations, where boss monsters and supporting entities (e.g., Guardians near the Elder Guardian, End Crystals for Ender Dragon healing) were spawned to ensure scenario complexity. During combat, real-time progress information (boss health, remaining End Crystals, etc.) was communicated through the chat system to support strategic planning. Task completion time was measured excluding initialization phases.
Evaluation metrics are defined as follows: health ratio = $\frac{1}{n}\sum_{i=1}^{n}\frac{h_{i}}{H_{\text{max}}}\times 100\%$ ($h_{i}$: agent’s remaining health, $H_{\text{max}}$: agent’s max health) and progress = $\frac{H_{\text{boss}}-h_{\text{boss}}}{H_{\text{boss}}}\times 100\%$ ($h_{\text{boss}}$: boss’s remaining health, $H_{\text{boss}}$: boss’s max health).
C.3 Adversarial PVP Task
@@
Similar to the Boss Combat Task, in this task, agents were provided with combat resources. Results were recorded when one team of agents was entirely defeated. After multiple experiments, victory rates and other relevant metrics were calculated. In order to promote fair confrontation, we give each team exactly the same initial supplies. Here is an example of initial inventory given to an agent : { ’wooden_sword’: 1, ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1, ’bow’: 1, ’arrow’: 200, ’golden_apple’: 20 }
C.4 Human-Agent Interaction
This section presents examples of human-agent interactions in dynamic environments, highlighting how agents and human players collaborate in various tasks.
Scenario 1: Human-Agent Collaboration in Resource Collection Tasks
Agent Steve’s Environment Information:
"time": "day",
"health": 20,
"hunger": 20,
"position": "(-223, 60, -159)",
"equipment": [
null,
null,
null,
null,
"crafting_table",
null
"inventory": {
"crafting_table": 1
"nearbyBlocks": [
"dirt",
"grass_block",
"oak_log",
"oak_leaves"
"nearbyEntities": "pig",
"nearbyPlayers": ["Alex", "Notch"]
Human Player’s Chat Message:
@@ C.4 Human-Agent Interaction
’Hey Steve, I’ve got some beef, can you hunt a pig for obtaining porkchops?’
Agent Steve’s Planning Action:
"skill": "obtainItem(bot, 10, \’porkchop\’)",
"interrupt": false,
@@
"reason": "Following the human players request to hunt a pig for obtaining porkchops. Since there are pigs nearby, I will proceed with hunting to gather porkchops."
Scenario 2: Human Player Directing Agents for Combat Tasks
Agent Steve’s Environment Information:
"time": "day",
"health": 1.2,
"hunger": 17,
"position": "(6, 66, 13)",
"equipment": [
"diamond_helmet",
"diamond_chestplate",
"diamond_leggings",
"diamond_boots",
"bow",
null
"inventory": {
"arrow": 189,
"golden_apple": 4,
"wooden_sword": 1,
"bow": 1
"nearbyBlocks": [
"end_stone",
"bedrock"
"nearbyEntities": "ender_dragon",
"nearbyPlayers": [
"Steve",
"Notch",
"Jeb"
Human Player’s Chat Message:
@@
’Steve, you are in grave danger. Cease your attack immediately and prepare to restore health!’
Agent Steve’s Planning Action:
"skill": "consumeItem(bot, \’golden_apple\’, false)",
"interrupt": true,
@@
"reason": "My health is critically low at 1.2, and I need to restore it immediately to continue fighting the ender_dragon effectively. I will follow the human players advice and prioritize my health."

### v2

@@ Appendix B Comprehensive Skill Library
Our proposed comprehensive skill library for Minecraft is implemented based on Mineflayer. To simulate the process of solving prerequisite steps in the actual game to eventually obtain specific items, this skill library has introduced a recursive task decomposition mechanism. From the implementation level, all interfaces can recursively invoke each other or themselves and pass parameters.
B.1. Minecraft Skill APIs
The Basic Skill Interfaces provide the foundational functions required to perform essential actions in the game, such as collecting items, crafting, mining, smelting, and combat. These functions are designed to be modular and reusable, allowing for seamless integration into more complex workflows.
Basic Skill Interfaces obtainItem(bot, count, type): Automates the collection of items. mineItem(bot, count, type, explore_direction, explore_time): Mines using specific tools. If the exploration parameters are not specified, the exploration direction is chosen randomly. If the item type is a subterranean mineral like diamond, the exploration direction is set to down (0, -1, 0). The exploration time limit defaults to five minutes (6000 ticks in Minecraft). craftItem(bot, count, type, need_crafting_table): Crafts items. If no parameter is specified, the crafting is done using a crafting table by default. smeltItem(bot, count, type, fuel): Smelts or cooks items in a furnace. If no parameters are provided, coal is used as the default fuel. collectItem(bot, count, type, function): Collects items by killing animals or using other special methods. chatMessage(bot, message, team_name): Sends a chat message to the team. getItemFromChest(bot, chest_position, items_to_get): Retrieves specific items from a chest at a given location. depositItemIntoChest(bot, chest_position, items_to_deposit): Deposits specific items into a chest at a given location. combatWithEntity(bot, mob_name, weapon, loop): Automatically equips the highest quality armor and weapon from the inventory to fight with an entity (e.g., animals or hostile mobs). If no parameters are specified, the default weapon is a sword. The loop parameter determines whether to continue fighting the same type of entity. combatWithPlayer(bot, player_name, weapon): Automatically equips the highest quality armor and weapon from the inventory to fight with a specific player. If no parameters are provided, the default weapon is a sword. initialInventory(bot, item_dict): Initializes the player’s inventory with specified items. equipBestToolOrArmor(bot, type): Automatically equips the highest quality tool, weapon, or armor of the specified type. listenChat(bot, player_name): Continuously listens to a player’s chat messages.
The Recursive-related Interfaces are designed to handle the dependencies between tasks. These functions determine the prerequisites for obtaining specific items, such as tools, materials, or entities, and ensure that all necessary steps are completed before proceeding with the main task.
All task dependencies are modeled as a Directed Acyclic Graph (DAG), where nodes represent tasks and edges represent prerequisite relationships. We have cataloged thousands of these intrinsic dependencies within Minecraft to support this recursive decomposition mechanism. This not only provides a methodological framework but also constitutes a significant engineering contribution by facilitating the automation of complex task sequences.
@@ B.1. Minecraft Skill APIs
Recursive-related Interfaces preTool(item): Retrieves the minimum prerequisite tool required for a specific item. preItem(item): Retrieves the prerequisite items required to craft a specific item and indicates whether a crafting table is necessary. preSmelt(item): Retrieves the prerequisite items required to smelt a specific item. preCollect(item): Retrieves the prerequisite entities required for a specific item. getFunc(item): Retrieves the method used to collect a specific item.
B.2. Implementation Details
Algorithm 2 demonstrates the implementation of the obtainItem function, which determines the collection method for a specific item and delegates the task to the corresponding sub-function (e.g., mining, crafting, smelting). This function serves as the entry point for recursive task decomposition.
Algorithm 2 obtainItem
1:function obtainItem($bot,cnt,type$)
2:  $func\leftarrow\textsc{getCollectionMethod}(type)$
3:  if $func=\text{``craft''}$ then
4:   return $\textsc{craftItem}(bot,cnt,type)$
5:  else if $func=\text{``mine''}$ then
6:   return $\textsc{mineItem}(bot,cnt,type)$
7:  else if $func=\text{``smelt''}$ then
8:   return $\textsc{smeltItem}(bot,cnt,type)$
9:  else
10:   return false
11:  end if
12:end function
The mineItem function shown in Algorithm 3 handles the process of mining blocks or ores using appropriate tools. If the required tool is not in the inventory, it recursively calls obtainItem to craft or gather the tool before proceeding with mining. This ensures that the bot is always equipped with the necessary tools for the task.
Algorithm 3 mineItem
1:function mineItem($bot,cnt,type,dir$)
2:  $tool\leftarrow\textsc{preTool}(type)$
3:  if $\textsc{invCnt}(tool)=0$ then
4:   $\textsc{obtainItem}(bot,1,tool)$
5:  end if
6:  $\textsc{equip}(bot,tool)$
7:  if $dir=\text{null}$ then
8:   $dir\leftarrow\textsc{randomDirection}$
9:  end if
10:  $\textsc{startMining}(bot,type,dir,time)$
11:  return $\textsc{verifyResult}(cnt,type)$
12:end function
The craftItem function shown in Algorithm 4 retrieves the recipe requirements for the specified item. If the required materials are insufficient, it recursively calls obtainItem to gather them before proceeding with crafting.
Algorithm 4 craftItem
1:function craftItem($bot,cnt,type$)
2:  $reqs\leftarrow\textsc{getRecipeReq}(type)$
3:  for (r_cnt,r_type)(r\_cnt,r\_type) in reqsreqs do
4:   if $\textsc{invCnt}(r\_type)<r\_cnt$ then
5:     $n\leftarrow r\_cnt-\textsc{invCnt}(r\_type)$
6:     $\textsc{obtainItem}(bot,n,r\_type)$
7:   end if
8:  end for
9:  $\textsc{executeCrafting}(type,cnt)$
10:  return $\textsc{verifyResult}(cnt,type)$
11:end function
The smeltItem function shown in Algorithm 5 checks if input items for smelting are available. If not, it recursively calls obtainItem to gather necessary materials before proceeding with smelting.
Algorithm 5 smeltItem
1:function smeltItem($bot,cnt,type$)
2:  $input\leftarrow\textsc{getSmeltingInput}(type)$
3:  if $\textsc{invCnt}(input)<cnt$ then
4:   $n\leftarrow cnt-\textsc{invCnt}(input)$
5:   $\textsc{obtainItem}(bot,n,input)$
6:  end if
7:  prepareFurnaceAndFuel
8:  $\textsc{executeSmelting}(type,cnt)$
9:  return $\textsc{verifyResult}(cnt,type)$
10:end function
@@ Appendix C Experiment Details
All experiments were conducted using the Qwen-Plus and Qwen-VL-plus model API provided by Alibaba Cloud. In the multiplayer local area network (LAN) server of Minecraft Java Edition version 1.19.4, agents connected to the game via different ports as an independent player and interacted with the environment and other agent players.
C.1. Resource Collection Task
Our framework supports tasks formulated as collecting any quantity of various types of items, represented by a requirement dictionary. This design ensures the richness and flexibility of supported tasks, enabling multi-agent collaboration to automatically collect resources based on given requirements. Such a setup makes our benchmark both general and scalable. However, due to inevitable time and API costs, we selected a set of representative resource collection tasks for our experiments. These tasks not only demonstrate the capabilities of our framework but also highlight common challenges faced by players in Minecraft, such as gathering essential survival tools (the Iron Tool Set task) or preparing for combat with powerful boss monsters (the Diamond Armor task). The following are the representative resource collection tasks we experimented with:
Task Definitions Iron Tool Set: { ’iron_pickaxe’: 1, ’iron_shovel’: 1, ’iron_hoe’: 1, ’iron_axe’: 1 } — A set of commonly used iron tools in Minecraft. Diamond Armor: { ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1 } — A full set of diamond armor in Minecraft. Redstone Devices: { ’repeater’: 1, ’piston’: 1, ’dropper’: 1 } — Common redstone components and devices in Minecraft. Navigation Kit: { ’compass’: 1, ’clock’: 1, ’map’: 1 } — A set of tools commonly used for navigation in Minecraft. Transport System: { ’minecart’: 1, ’rail’: 16, ’powered_rail’: 6 } — Minecarts and tracks used to complete a transportation system. Food Supplies: { ’beef’: 1, ’chicken’: 1, ’porkchop’: 1 } — Common animal meat food items in Minecraft. Building Materials: { ’stone_bricks’: 4, ’glass’: 4, ’iron_door’: 1 } — Building materials commonly used for constructing doors, windows, and walls. Storage System: { ’hopper’: 1, ’chest’: 1, ’barrel’: 1 } — Tools commonly used for storing items in Minecraft.
@@ C.1. Resource Collection Task
C.2. Boss Combat Task
In Minecraft, there are three primary world dimensions: Overworld, Nether, and End. Each dimension hosts powerful boss monsters whose defeat is considered the pinnacle challenge for players of Minecraft. Based on this, we have predefined three combat tasks:
  • Elder Guardian (Overworld): This task involves defeating the formidable boss Elder Guardian within complex underwater terrain of an ocean monument. Agents must also contend with other surrounding monsters, adding complexity that requires the MAS to develop effective combat strategies.
  • Wither (Nether): This task involves battling the powerful boss Wither in the Nether. Agents face additional challenges from smaller enemies such as piglins and ghasts, amidst a landscape filled with lava and other hazards. When the Wither’s health drops below half, it enters a berserk state where it becomes immune to ranged attacks, necessitating the MAS to dynamically adjust combat strategies based on real-time observations.
  • Ender Dragon (End): Defeating the Ender Dragon in the End is considered the ultimate challenge in Minecraft. Multiple agents need to strategically cooperate by first destroying the end crystals located atop obsidian pillars to disable the dragon’s healing mechanism before engaging in battle with the dragon and its surrounding endermen, requiring the MAS to adjust tactics adaptively.
@@ C.2. Boss Combat Task
Prior to each combat, agents were equipped with standardized combat resources, including a full set of diamond armor (validated as efficiently collectible through our resource collection experiments), bow and arrows, and some consumables. In all experiments, the standardized combat resources remained consistent. A single LLM call initialized resource allocation, distributing equipment to each agent through a dictionary-based assignment system. Agents were then teleported to designated combat locations, where boss monsters and supporting entities (e.g., Guardians near the Elder Guardian, End Crystals for Ender Dragon healing) were spawned to ensure scenario complexity. During combat, real-time progress information (boss health, remaining End Crystals, etc.) was communicated through the chat system to support strategic planning. Task completion time was measured excluding initialization phases.
Evaluation metrics are defined as follows:
$\text{health ratio}=\frac{1}{n}\sum_{i=1}^{n}\frac{h_{i}}{H_{\text{max}}}\times 100\%$ (10)
where $h_{i}$ is the remaining health of agent $i$, and $H_{\text{max}}$ is the maximum health of the agent.
$\text{progress}=\frac{H_{\text{boss}}-h_{\text{boss}}}{H_{\text{boss}}}\times 100\%$ (11)
where $h_{\text{boss}}$ is the remaining health of the boss monster, and $H_{\text{boss}}$ is the maximum health of the boss monster.
C.3. Adversarial PVP Task
@@
In addition to ablation studies, we design a task scenario where two frameworks can be directly compared, namely Adversarial PVP Task. Similar to the Boss Combat Task, in this task, agents were provided with combat resources. Then, the two teams of agents engaged in battles with each other. Results were recorded when one team of agents was entirely defeated. After multiple experiments, victory rates and other relevant metrics were calculated. In order to promote fair confrontation, we give each team exactly the same initial supplies. Here is an example of initial inventory given to an agent : { ’wooden_sword’: 1, ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1, ’bow’: 1, ’arrow’: 200, ’golden_apple’: 20 }
C.4. Ablation Study
Our ablation studies respectively focused on the comprehensive skill library and its recursive task decomposition, the centralized memory system and the parallelized planning-acting mechanism, which are three core contributions proposed in our paper. Although it is challenging to conduct fair comparisons between different approaches and frameworks in agent-based research due to varying architectures and objectives, each ablation here can be regarded as a methodological baseline. For example, ablating the parallelized planning-acting mechanism is actually taking the traditional serialized planning-acting architecture as the baseline.
C.5. Scale Robustness Analysis
Our method theoretically supports multi-agent systems of any scale. Regardless of the scale, each agent is independently equipped with parallel planning threads and acting threads and interacts with the environment in real time. As for the acting thread of a single agent, its actions are completely unaffected by the number of agents in the entire system. However, for the planning thread, under our settings, the length of the chat log window input for each LLM call is equal to the number of system agents, which means that as the system scale increases, the number of input tokens for LLM inference continues to increase. So we conducted an experiment on this. By adjusting the window length of the chat log, we sampled different numbers of chat entries for LLM planning and counted their inference duration and the number of input and output tokens.
As shown in the main paper, experiment results show that the LLM inference time tends to stabilize rather than grow continuously with increasing agents and the total token cost grows approximately linearly with the number of agents. This may be attributed to the fact that we utilize the Qwen-Plus API, which likely incorporates advanced optimization techniques such as vLLM acceleration, enabling the system to efficiently handle increased tokens without significantly impacting inference latency, thereby maintaining stable performance even as the input size grows. The input tokens increases linearly with the number of agents, as more chat entries are processed during each planning step. However, the output tokens remain stable due to the constraints imposed by the prompt format, which limits the length of generated responses. Overall, our method demonstrates robustness to the scale of multi-agent systems.
Appendix D Human-Agent Interaction
Our framework also supports flexible human-agent interaction, allowing human players to either actively participate in task execution alongside agents, facilitating dynamic cooperation within the game environment, or to take on a guiding role by providing instructions to agents for task division, coordination, and strategic decision-making. This section presents examples of human-agent interactions in Minecraft, highlighting how AI agents and human players collaborate in various tasks.
D.0.1. Scenario 1: Human-Agent Collaboration
Agent Steve’s Environment Information { time”: day”, health”: 20, hunger”: 20, position”: ”(-223, 60, -159)”, equipment”: [ null, null, null, null, crafting_table”, null ], inventory”: { crafting_table”: 1 }, nearbyBlocks”: [ dirt”, grass_block”, oak_log”, oak_leaves ], nearbyEntities”: pig”, nearbyPlayers”: [”Alex”, Notch”] }
@@ D.0.1. Scenario 1: Human-Agent Collaboration
Human Player’s Chat Message ’Hey Steve, I’ve got some beef, can you hunt a pig for obtaining porkchops?’
@@
Agent Steve’s Planning Action { skill”: obtainItem(bot, 10, \’porkchop\’)”, interrupt”: false, reason”: Following the human players request to hunt a pig for obtaining porkchops. Since there are pigs nearby, I will proceed with hunting to gather porkchops.” }
Scenario 2: Human Player Directing Agents
Agent Steve’s Environment Information { time”: day”, health”: 1.2, hunger”: 17, position”: ”(6, 66, 13)”, equipment”: [ diamond_helmet”, diamond_chestplate”, diamond_leggings”, diamond_boots”, bow”, null ], inventory”: { arrow”: 189, golden_apple”: 4, bow”: 1 }, nearbyBlocks”: [ end_stone ], nearbyEntities”: ender_dragon”, nearbyPlayers”: [ Jeb ] }
@@
Human Player’s Chat Message ’Steve, you’re in grave danger. Cease attack immediately and prepare to restore health!’
@@
Agent Steve’s Planning Action { skill”: consumeItem(bot, \’golden_apple\’, false)”, interrupt”: true, reason”: My health is critically low at 1.2, and I need to restore it immediately to continue fighting the ender_dragon effectively. I will follow the human players advice and prioritize my health.” }